OpenStudy (anonymous):

If r1 and r2 are the roots of the quadratic equation ax^2+bx+c=0, show that r1+r2=-b/a and r1r2=c/a

OpenStudy (anonymous):

If r1 and r2 are roots then (x-r1)(x-r2)=0

OpenStudy (anonymous):

x^2-x(r2)-x(r2)+(r1*r2)=x^2-x(r1+r2)+(r1^2*r2^2) Your quadratic equation is y=x^2-x(r1+r2)+(r1*r2) compare to y=ax^2+bx+c

OpenStudy (anonymous):

a=1 , b=-(r1+r2) , and c=r1*r2

OpenStudy (anonymous):

but a =1, -b=-b/a=(r1+r2) c=c/a=r1*r2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

thank you!